# TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision

### v1

@@ Abstract.
Hierarchical text classification aims to categorize each document into a set of classes in a label taxonomy. Most earlier works focus on fully or semi-supervised methods that require a large amount of human annotated data which is costly and time-consuming to acquire. To alleviate human efforts, in this paper, we work on hierarchical text classification with the minimal amount of supervision: using the sole class name of each node as the only supervision. Recently, large language models (LLM) show competitive performance on various tasks through zero-shot prompting, but this method performs poorly in the hierarchical setting, because it is ineffective to include the large and structured label space in a prompt. On the other hand, previous weakly-supervised hierarchical text classification methods only utilize the raw taxonomy skeleton and ignore the rich information hidden in the text corpus that can serve as additional class-indicative features. To tackle the above challenges, we propose TELEClass, Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification, which (1) automatically enriches the label taxonomy with class-indicative topical terms mined from the corpus to facilitate classifier training and (2) utilizes LLMs for both data annotation and creation tailored for the hierarchical label space. Experiments show that TELEClass can outperform previous weakly-supervised hierarchical text classification methods and LLM-based zero-shot prompting methods on two public datasets.
Weakly-Supervised Text Classification, Hierarchical Text Classification, Taxonomy Enrichment, Large Language Model
$*{}^{*}$qual Contribution.
1. Introduction
@@
Hierarchical text classification, aiming to classify documents into one or multiple classes in a label taxonomy, is an essential and fundamental task in text mining and natural language processing. Compared with standard text classification where label space is flat and relatively small (e.g., less than 20 classes), hierarchical text classification is more challenging given the larger and more structured label space and the existence of fine-grained and long-tail classes in the taxonomy. Most earlier works tackle this task in fully supervised (You et al., 2019; Huang et al., 2019; Chen et al., 2021) or semi-supervised settings (Berthelot et al., 2019; Gururangan et al., 2019), and different models are proposed to learn from a substantial amount of human-labeled data. However, acquiring human annotation is often costly, time-consuming, and not scalable.
@@
Figure 1. An example document tagged with 3 classes. We automatically enrich each node with class-indicative terms and utilize LLMs to facilitate classification.
@@
Recently, large language models (LLM) such as GPT-4 (OpenAI, 2023) and Llama2 (Touvron et al., 2023) have demonstrated strong performance in flat text classification through zero-shot or few-shot prompting (Sun et al., 2023). However, applying LLMs in hierarchical settings, with large and structured label spaces, remains challenging. Directly including hundreds of classes in prompts is ineffective, leading to structural information loss, increased computational costs, and diminished clarity for LLMs in focusing on critical information. Along another line of research, Meng et al. (2019) propose to train a moderate-size text classifier by utilizing a small set of keywords or labeled documents for each class. However, compiling keyword lists for hundreds of classes and obtaining representative documents for each specific and niche category still demand significant human efforts. Shen et al. (2021) study the hierarchical text classification with minimal supervision, which takes the class name as the only supervision signal. Specifically, they introduce TaxoClass, a method that generates pseudo-labels with a textual entailment model for classifier training. However, this method overlooks additional class-relevant features in the text and suffers from unreliable pseudo-label selection due to the entailment model’s domain shift.
@@
In this study, we advance minimally supervised hierarchical text classification by integrating corpus-based taxonomy enrichment and leveraging LLMs tailored for the hierarchical label structure. First, we enrich the existing label taxonomy skeleton with automatically extracted class-specific topical terms from the corpus, enhancing supervision signals and thus improving the pseudo-label quality for classifier training. Second, we explore the application of LLMs in hierarchical text classification from two perspectives. Specifically, we enhance LLM annotation efficiency and effectiveness through a taxonomy-guided candidate search and optimize LLM-generated document accuracy by using taxonomy paths to create more precise pseudo-data.
@@
Leveraging the above ideas, we introduce TELEClass: Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification. TELEClass consists of four major steps: (1) LLM-Enhanced Core Class Annotation, where we identify document “core classes” (i.e., fine-grained classes that most accurately describe the documents) by first using a textual entailment model for top-down candidate search on the taxonomy and then prompting an LLM to select the most precise core classes. (2) Corpus-Based Taxonomy Enrichment, where we analyze the taxonomy structure and the roughly classified documents to identify class-indicative topical terms through semantic and statistical analysis, enriching the label taxonomy with more features. For example, the “conditioner” class in Figure 1 is enriched with key terms like “moisture” and “soft hair”, which distinguish it from its sibling classes. (3) Core Class Refinement with Enriched Taxonomy, where we embed documents and classes based on the enriched label taxonomy and refine the initially selected core classes by identifying the most similar classes for each document. (4) Text Classifier Training with Path-Based Data Augmentation, where we sample label paths from the taxonomy and guide the LLM to generate pseudo documents most accurately describing these fine-grained classes. Finally, we train the text classifier on two types of pseudo-labels, the core classes and the generated data, with a simple text-matching network and multi-label training strategy.
@@
The contributions of this paper are summarized as follows: (1) We propose TELEClass, a new method for minimally supervised hierarchical text classification, which requires only the class names of the label taxonomy as supervision to train a multi-label text classifier. (2) We propose to enrich the label taxonomy with class-indicative topical terms mined from the text corpus, based on which we utilize an embedding-based document-class matching method to improve the pseudo-label quality. (3) We study two ways of adopting large language models to the hierarchical text classification setting, which can improve the pseudo-label quality and solve the data scarcity issue for fine-grained classes. (4) Extensive experiments on two datasets show that TELEClass can outperform strong weakly-supervised hierarchical text classification methods and zero-shot LLM prompting methods.  [^1] Code and datasets are available upon request now and will later be available online.
2. Problem Definition
@@
Following Shen et al. (2021), we define the minimally-supervised hierarchical text classification task as to train a text classifier that can categorize each document into multiple nodes on a label taxonomy by using the name of each node as the only supervision. For example, in Figure 1, the input document is classified as “hair care”, “shampoo”, and “scalp treatment”.
@@
3. Methodology
@@
Figure 2. Overview of the TELEClass framework.
3.1. LLM-enhanced Core Class Annotation
@@
Inspired by previous studies, we first tag each document with its “core classes”, which are defined as a set of classes that can describe the document most accurately (Shen et al., 2021). This process also mimics the process of human performing hierarchical text classification: first, select a set of most essential classes for the document and then trace back to their ancestors to complete the labeling. For example, in Figure 1, by first tagging the document with “shampoo” and “scalp treatment”, we can easily find its complete set of classes.
To mine the core classes for each document, existing methods utilize a textual entailment model (e.g., RoBERTa-large-MNLI) and combine tree-based top-down candidate searching with corpus-level statistical analysis for core class selection. Specifically, the similarity score between a document and a class is defined as the entailment score from the document to a class-specific hypothesis such as “This example is [Class Name]”. These similarity scores facilitate a top-down search in the taxonomy to shortlist candidate classes, which are then evaluated across the corpus to determine the most reliable core classes.
However, there are two limitations of this method. First, the textual entailment model, not being directly trained for document classification, offers inconsistent zero-shot performance, particularly in distinguishing closely related, fine-grained classes within a hierarchical structure. Second, given a pair of premise (document) and hypothesis, the entailment model is trained to predict a distribution over three classes, namely entailment, neutral, and contradiction. Therefore, it is questionable whether the generated scores can support corpus-level comparison, because the scores between different pairs may not be directly comparable. Both limitations will lead to noisy core class selection and adversely impact the classifier performance.
In this work, we propose to utilize an LLM to enhance the core class annotation step. These LLMs are trained on very large textual data and human preference data and shown to have superior zero-shot ability for the document classification task with a small and flat label space (Sun et al., 2023). However, LLMs cannot effectively understand the hierarchical label space, which is large and structured and cannot be easily represented in a prompt. Therefore, we precede the LLM annotation with the top-down candidate search method using a textual entailment model.
Specifically, for each document $d_{i}$, we start from the Root node at level 0 of the taxonomy by adding it into a queue. For each class at level $l$ in the queue, we select $l+3$ of its children that have the highest similarity to document $d_{i}$, where the similarity score $sim^{\text{entail}}(c,d_{i})$ is obtained from a textual entailment model. Then, among all the selected children at level $l+1$, we keep the top $(l+2)^{2}$ of them with the highest path-based similarity scores and add to the queue, where the path-based similarity $sim^{\text{path}}(c,d_{i})$ is defined recursively as,
$sim^{\text{path}}(\text{Root},d)$ $=1,$
$sim^{\text{path}}(c,d)$ $=\max_{c_{p}\in Par(c)}sim^{\text{path}}(c_{p},d_{i})\cdot sim^{\text{entail}}(c,d_{i}),$
where $Par(c)$ is the set of all parents of $c$. This process finishes when there is no class in the queue, and all the nodes (other than Root) ever pushed into the queue are treated as candidate core classes for document $d_{i}$. To use an LLM to annotate core classes, we first construct an instruction for each document, which asks the LLM to select the most accurate classes for the document from its candidate core classes. Then, the document is fed to the LLM as a query and it will generate a list of core classes according to the document and candidate core classes.
Our LLM-enhanced core class annotation step will assign an initial set of core classes (denoted as $\mathbb{C}^{0}_{i}$) for each document $d_{i}\in\mathcal{D}$. In the next two steps, we introduce how TELEClass leverages the initial core classes to enrich the label taxonomy with class-indicative features. This enriched taxonomy is subsequently employed to refine the initial core classes, utilizing embedding-based similarity scores that are more comparable across documents.
3.2. Corpus-Based Taxonomy Enrichment
Due to the hierarchical nature of taxonomy, fined-grained classes at lower levels are often hard to distinguish with only class label names. Therefore, we propose to enrich the weak supervision, i.e. the label taxonomy, with a set of class-indicative topical terms for each class. These enriched terms can help better interpret the class and thus facilitate text classification. For example, in Figure 2, “shampoo” and “conditioner” are two fine-grained classes that are similar to each other. We can effectively separate the two classes by identifying a set of class-specific terms such as “flakes” for “shampoo” and “moisture” for “conditioner”.
In this work, we combine statistical document occurrence features and semantic similarity and further contrast a class with its siblings to identify a set of key terms that can distinguish it from its most similar nodes in the taxonomy. Notice that, because the input taxonomy $\mathcal{T}$ is a directed acyclic graph rather than a simple tree, a node $c$ can have multiple sets of siblings, each of which corresponds to one of $c$’s parent nodes. Besides, because we have already tagged each document with its core classes, we can use them as rough classification results and form a document cluster for each class to help the enrichment step.
Formally, given a class $c\in\mathcal{C}$ and its siblings corresponding to one of its parents $c_{p}$, $Sib(c,c_{p})=\{c^{\prime}\in\mathcal{C}|\langle c_{p},c^{\prime}\rangle\in\mathcal{R}\}$, $\mathbf{c}_{p}\in Par(c)$, we find a set of class-indicative topical terms of $c$ corresponding to $c_{p}$, denoted as $T(c,c_{p})=\{t_{1},t_{2},\dots,t_{k}\}$. Each term in $T(c,c_{p})$ can signify the class $c$ and distinguish it from its siblings under $c_{p}$. To identify these class-indicative terms, we construct a set of documents $D^{0}_{c}$ for each class $c$, containing all the documents whose initial core classes contain $c$ or its descendants,
(1) $\small D^{0}_{c}=\left\{d_{i}\in D|\exists c^{\prime}\in\mathbb{C}^{0}_{i},c^{\prime}\in\{c\}\cup Des(c)\right\},$
where $Des(c)$ is the set of all descedants of $c$. We consider the following 3 factors for class-indicative term selection: popularity, distinctiveness, and semantic similarity.
Popularity. First, a class-indicative term of $c$ should be frequently mentioned by the documents about $c$, denoted as its popularity within $c$. We calculate the popularity of a term $t$ within a class $c$ as,
(2) $\small pop(t,c)=\log(1+df(t,D^{0}_{c})),$
where the document frequency $df(t,D)$ stands for the number of documents in $D$ that mention $t$.
Distinctiveness. Second, a class-indicative term $t$ for a class $c$ should also be infrequent in its siblings so that a document containing $t$ likely indicates that it is about $c$ instead of $c$’s siblings. We define the distinctiveness score of a term $t$ for a class $c$, corresponding to its siblings under $c_{p}$, as,
(3) $\small dist(t,c,c_{p})=\frac{\exp(BM25(t,D^{0}_{c}))}{1+\sum_{c^{\prime}\in Sib(c,c_{p})}\exp(BM25(t,D^{0}_{c^{\prime}}))},$
where $BM25(\cdot,\cdot)$ denotes the BM25 relevance function (Robertson and Zaragoza, 2009).
[…]
Finally, we define the affinity score between a term $t$ and a class $c$ corresponding to parent $p$ to be the geometric mean of the above scores, with a weight parameter $\alpha$,
(4) $\small aff(t,c,c_{p})=pop(t,c)^{\alpha}\cdot dist(t,c,c_{p})^{1-\alpha}\cdot\cos(\mathbf{t},\mathbf{c}).$
@@ 3.2. Corpus-Based Taxonomy Enrichment
Taxonomy Enrichment. We first apply a phrase mining tool, AutoPhrase (Shang et al., 2018), to mine quality single-token and multi-token phrases from the corpus to serve as candidates for taxonomy enrichment. Then, for each class $c$ and each of its parents $p$, we select the top-$k$ terms with the highest affinity scores with $c$ corresponding to $p$, denoted as $T(c,c_{p})$. Then, we aggregate the terms corresponding to each parent of $c$ to get the final enriched class-indicative topical terms for class $c$ as follows
(5) $\small T_{c}=\bigcup_{c_{p}\in Par(c)}T(c,c_{p}).$
3.3. Core Class Refinement with Enriched Taxonomy
With the enriched class-indicative terms for each class, we propose to further utilize them to refine the initial core classes. In this paper, we adopt an embedding-based document-class matching method, which has two distinct advantages in the hierarchical setting. (1) By pre-computing the document and class representations, calculating their similarity is much more efficient than the textual entailment-based method, promoting a more comprehensive score comparison on the entire taxonomy instead of just on a sub-tree. (2) While textual entailment model is trained to predict an entailment score for each pair of document and class, which follows an unknown distribution and is unreliable for corpus-level comparison, the embedding similarity is more general and easier to compare across classes and documents.

### v2

@@ Abstract
Hierarchical text classification aims to categorize each document into a set of classes in a label taxonomy. Most earlier works focus on fully or semi-supervised methods that require a large amount of human annotated data which is costly and time-consuming to acquire. To alleviate human efforts, in this paper, we work on hierarchical text classification with the minimal amount of supervision: using the sole class name of each node as the only supervision. Recently, large language models (LLM) show competitive performance on various tasks through zero-shot prompting, but this method performs poorly in the hierarchical setting, because it is ineffective to include the large and structured label space in a prompt. On the other hand, previous weakly-supervised hierarchical text classification methods only utilize the raw taxonomy skeleton and ignore the rich information hidden in the text corpus that can serve as additional class-indicative features. To tackle the above challenges, we propose TELEClass, Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification, which (1) automatically enriches the label taxonomy with class-indicative terms to facilitate classifier training and (2) utilizes LLMs for both data annotation and creation tailored for the hierarchical label space. Experiments show that TELEClass can outperform previous weakly-supervised methods and LLM-based zero-shot prompting methods on two public datasets.
TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision
Yunyi Zhang1, Ruozhen Yangthanks: Equal contribution1, Xueqiang Xu11footnotemark: 11, Rui Li11footnotemark: 12, Jinfeng Xiao1, Jiaming Shen3, Jiawei Han1, 1University of Illinois Urbana-Champaign, 2University of Science and Technology of China, 3Google DeepMind, yzhan238@illinois.edu
1 Introduction
@@
Hierarchical text classification, aiming to classify documents into one or multiple classes in a label taxonomy, is an essential and fundamental task in text mining and natural language processing. Compared with standard text classification where label space is flat and relatively small (e.g., less than 20 classes), hierarchical text classification is more challenging given the larger and more structured label space and the existence of fine-grained and long-tail classes. Most earlier works tackle this task in fully supervised You et al. (2019); Huang et al. (2019); Chen et al. (2021) or semi-supervised settings Berthelot et al. (2019); Gururangan et al. (2019), and different models are proposed to learn from a substantial amount of human-labeled data. However, acquiring human annotation is often costly, time-consuming, and not scalable.
@@
Figure 1: An example document tagged with 3 classes. We automatically enrich each node with class-indicative terms and utilize LLMs to facilitate classification.
@@
Recently, large language models (LLM) such as GPT-4 OpenAI (2023) and Claude 3 Anthropic (2024) have demonstrated strong performance in flat text classification through zero-shot or few-shot prompting Sun et al. (2023). However, applying LLMs in hierarchical settings, with large and structured label spaces, remains challenging. Directly including hundreds of classes in prompts is ineffective, leading to structural information loss, increased computational costs, and diminished clarity for LLMs at distinguishing class-specific information. Along another line of research, Meng et al. (2019) propose to train a moderate-size text classifier by utilizing a small set of keywords or labeled documents for each class. However, compiling keyword lists for hundreds of classes and obtaining representative documents for each specific and niche category still demand significant human efforts. Shen et al. (2021) study the hierarchical text classification with minimal supervision, which takes the class name as the only supervision signal. Specifically, they introduce TaxoClass which generates pseudo labels with a textual entailment model for classifier training. However, this method overlooks additional class-relevant features in the text and suffers from the unreliable pseudo label selection due to domain shift of the entailment model.
@@
In this study, we advance minimally supervised hierarchical text classification by enriching the weak supervision with additional textual clues and leveraging LLMs tailored for the hierarchical label structure. First, we enrich the existing label taxonomy with class-specific terms derived from two sources: LLM generation and automated extraction from the corpus. For example, the “conditioner” class in Figure 1 is enriched with key terms like “moisture” and “soft hair”, which distinguish it from other classes. These terms enhance the supervision signal by combining knowledge from LLMs and text corpus and improve the pseudo label quality for classifier training. Second, we further explore the application of LLMs in hierarchical text classification from two perspectives. Specifically, we enhance LLM annotation efficiency and effectiveness through a taxonomy-guided candidate search and optimize LLM document generation to create more precise pseudo data by using taxonomy paths.
@@
Leveraging the above ideas, we introduce TELEClass: Taxonomy Enrichment and LLM-Enhanced weakly-supervised hierarchical text Classification. TELEClass consists of four major steps: (1) LLM-Enhanced Core Class Annotation, where we identify document “core classes” (i.e., fine-grained classes that most accurately describe the documents) by first enriching the taxonomy with LLM-generated key terms and then finding candidate classes with a top-down tree search algorithm for LLM to select the most precise core classes. (2) Corpus-Based Taxonomy Enrichment, where we analyze the taxonomy structure to additionally identify class-indicative topical terms through semantic and statistical analysis on the corpus. (3) Core Class Refinement with Enriched Taxonomy, where we embed documents and classes based on the enriched label taxonomy and refine the initially selected core classes by identifying the most similar classes for each document. (4) Text Classifier Training with Path-Based Data Augmentation, where we sample label paths from the taxonomy and guide the LLM to generate pseudo documents most accurately describing these fine-grained classes. Finally, we train the text classifier on two types of pseudo labels, the core classes and the generated data, with a simple text matching network and multi-label training strategy.
@@
The contributions of this paper are summarized as follows: (1) We propose TELEClass, a new method for minimally supervised hierarchical text classification, which requires only the class names of the label taxonomy as supervision to train a multi-label text classifier. (2) We propose to enrich the label taxonomy with class-indicative terms, based on which we utilize an embedding-based document-class matching method to improve the pseudo label quality. (3) We study two ways of adopting large language models to help hierarchical text classifier training, which can improve the pseudo label quality and solve the data scarcity issue for fine-grained classes. (4) Extensive experiments on two datasets show that TELEClass can outperform strong weakly-supervised hierarchical text classification methods and zero-shot LLM prompting methods.
2 Problem Definition
@@
The minimally-supervised hierarchical text classification task aims to train a text classifier that can categorize each document into multiple nodes on a label taxonomy by using the name of each node as the only supervision Shen et al. (2021). For example, in Figure 1, the input document is classified as “hair care”, “shampoo”, and “scalp treatment”.
@@
Figure 2: Overview of the TELEClass framework.
We assume the label taxonomy to be a DAG instead of a tree, because it aligns better with real applications as one node can have multiple parents with different meanings. It is also more challenging because the classifier needs to assign a document multiple labels in different levels and paths.
3 Methodology
@@
3.1 LLM-Enhanced Core Class Annotation
@@
Inspired by previous studies, we first tag each document with its “core classes”, which are defined as a set of classes that can describe the document most accurately Shen et al. (2021). This process also mimics the process of human performing hierarchical text classification: first select a set of most essential classes for the document and then trace back to their relevant classes to complete the labeling. For example, in Figure 1, by first tagging the document with “shampoo” and “scalp treatment”, we can easily find its complete set of classes.
In this work, we propose to enhance the core class annotation process of previous methods with the power of LLMs. To utilize LLMs for core class annotation, we apply a structure-aware candidate core class selection method to reduce the label space for each document. This step is necessary because LLMs can hardly comprehend a large, structured hierarchical label space that cannot be easily represented in a prompt. We first define a similarity score between a class and a document that will be used in candidate selection. To better capture the semantics, we propose to use LLMs to generate a set of class relevant keywords to enrich the raw taxonomy structure and consolidate the meaning of each class. For example, “shampoo” and “conditioner” are two fine-grained classes that are similar to each other. We can effectively separate the two classes by identifying a set of class-specific terms such as “flakes” for “shampoo” and “moisture” for “conditioner”. We prompt an LLM to enrich the raw label taxonomy with a set of key terms for each class, denoted as $T^{\text{LLM}}_{c}$ for the class $c$. To ensure the generated terms can uniquely identify $c$, we ask the LLM to generate terms that are relevant to $c$ and its parent while irrelevant to siblings of $c$. With this enriched set of terms for each class, we define the similarity score between a document $d$ and a class $c$ as the maximum cosine similarity with the key terms:
$sim(c,d)=\max_{t\in T^{\text{LLM}}_{c}}\cos(\vec{t},\vec{d}),\vspace{-0.5em}$ (1)
where $\vec{\circ}$ denotes a vector representation by a pre-trained semantic encoder (e.g., Sentence Transformer Reimers and Gurevych (2019)).
This newly defined similarity measure is then used for candidate core class selection. Given a document, we start from the root node at level $l=0$, select the $l+3$ most similar children classes to the document at level $l$ using the similarity score defined above, and continue to the next level with only the selected classes. The increasing number of selected nodes accounts for the growing number of classes when going deeper into the taxonomy. Finally, all the classes ever selected in this process will be the candidate core classes for this document, which share the most similarity with the document according to the label hierarchy.  [^1] Refer to Shen et al. (2021) for more details on the tree search algorithm.
Finally, we instruct an LLM to select the core classes for each document from the selected candidates, which produces an initial set of core classes (denoted as $\mathbb{C}^{0}_{i}$) for each document $d_{i}\in\mathcal{D}$.
3.2 Corpus-Based Taxonomy Enrichment
In the previous step, we enrich the raw taxonomy structure with LLM-generated key terms, which are derived from the general knowledge of LLMs but may not accurately reflect the corpus-specific knowledge. Therefore, we propose further enriching the classes with class-indicative terms mined from the text corpus. By doing this, we can combine the general knowledge of LLMs and corpus-specific knowledge to better enhance the very weak supervision, which is essential for correctly understanding fine-grained classes that are hard to distinguish. Formally, given a class $c\in\mathcal{C}$ and its siblings corresponding to one of its parents $c_{p}$, $Sib(c,c_{p})=\{c^{\prime}\in\mathcal{C}|\langle c_{p},c^{\prime}\rangle\in\mathcal{R}\}$, $\mathbf{c}_{p}\in Par(c)$, we find a set of corpus-based class-indicative terms of $c$ corresponding to $c_{p}$, denoted as $T(c,c_{p})=\{t_{1},t_{2},\dots,t_{k}\}$. Each term in $T(c,c_{p})$ can signify the class $c$ and distinguish it from its siblings under $c_{p}$.
We first collect a set of relevant documents $D^{0}_{c}\subset\mathcal{D}$ for each class $c$, which contains all the documents whose initial core classes contain $c$ or its descendants. Then, inspired by Tao et al. (2018); Zhang et al. (2023b), we consider the following three factors for class-indicative term selection: (1) Popularity: a class-indicative term $t$ of a class $c$ should be frequently mentioned by its relevant documents, which is quantified by the document frequency $\log(1+df(t,D^{0}_{c}))$. (2) Distinctiveness: a class-indicative term $t$ for a class $c$ should be infrequent in its siblings, which is quantified as the softmax of $BM25(t,D^{0}_{c})$ relevance scores over the set of siblings, $Sib(c,c_{p})$. (3) Semantic similarity: a class-indicative term should also be semantically similar to the class name, which is quantified as the cosine similarity between their embeddings derived from a pre-trained encoder (e.g., BERT Devlin et al. (2019)). Finally, we define the affinity score between a term $t$ and a class $c$ corresponding to parent $p$ to be the geometric mean of the above scores, denoted as $\text{aff}(t,c,c_{p})$.
@@ 3.2 Corpus-Based Taxonomy Enrichment
To enrich the taxonomy, we first apply a phrase mining tool, AutoPhrase Shang et al. (2018), to mine quality single-token and multi-token phrases from the corpus as candidate terms [^2] Our method is flexible with any kinds of phrase mining methods like Gu et al. (2021).. Then, for each class $c$ and each of its parents $c_{p}$, we select the top-$k$ terms with the highest affinity scores with $c$ corresponding to $p$, denoted as $T(c,c_{p})$. Then, we take the union of these corpus-based terms together with the LLM-generated terms in the previous step to get the final enriched class-indicative terms for class $c$, $T_{c}=\left(\bigcup_{c_{p}\in Par(c)}T(c,c_{p})\right)\bigcup T^{\text{LLM}}_{c}$.
3.3 Core Class Refinement with Enriched Taxonomy
With the enriched class-indicative terms for each class, we propose to further utilize them to refine the initial core classes. In this paper, we adopt an embedding-based document-class matching method. Unlike previous methods in flat text classification Wang et al. (2021a) that use keyword-level embeddings to estimate document and class representations, here, we are able to define class representations directly based on document-level embeddings thanks to the rough class assignments we created in the core class annotation step (c.f. Sec. 3.1).